OpenStudy (anonymous):

A biologist is studying growth in space. He wants to simulate Earth's gravitational field, so he positions the plants on a rotating platform in the spaceship. The distance of each plant from the central axis of rotation is r = 0.37 m. What angular speed is required?

OpenStudy (theeric):

Hi! Do you have an attempt at a solution?

OpenStudy (anonymous):

not really since i dont even know where to start. they only gave me the radius. dont i need more information?

OpenStudy (theeric):

Nope! Do you know about centripetal force or centrifugal force?

OpenStudy (anonymous):

does this have to do with omega? T=2pi/w

OpenStudy (theeric):

Sort of! It's actually asking for \(\omega\)! That is the angular velocity! The problem concerns centrifugal force. When you're on the edge of something spinning, you feel like something's pushing you away from the center of your revolution. Like, if you grab a friend's hand and you both start rotating, you feel it get harder to hold on!

OpenStudy (theeric):

That, in space, is going to simulate gravitational force for that plant.

OpenStudy (anonymous):

ok so is angular speed the same thing as angular velocity?

OpenStudy (anonymous):

so omega=2pi/T but i dont have time and im given the radius instead. what should i don in this case?

OpenStudy (theeric):

Well, speed doesn't consider direction, and velocity does, but that's the only difference. They are measured in either degrees per second, radians per second, radians per minute, or things like that.

OpenStudy (theeric):

Well, you're given the radius for a specific reason. The centrifugal force that will be like gravitational force depends on \(\omega\) and the radius \(r\).

OpenStudy (theeric):

What you want is centripetal acceleration to equal what gravitational acceleration would be on Earth. Centripetal acceleration is \(a_c=\dfrac{v^2}{r}\) normally. But we can find it for \(\omega\).

OpenStudy (anonymous):

ah ok.

OpenStudy (theeric):

\(\omega=\dfrac{v}{r}\) So what is \(v\)?

OpenStudy (anonymous):

wouldn't that be the pull of gravity? 9.8m/s

OpenStudy (anonymous):

w=26.5 rads/s

OpenStudy (theeric):

\(v\)? No, sorry! That is the tangential velocity - tangent to your path. Like, if you're driving in a circle, it's what the car's speedometer would say.

OpenStudy (theeric):

That's not what I got, but I see what you did. We don't know \(v\). We want \(a_c=\dfrac{v^2}{r}\), so we need to solve for \(v\). But hey, it's related to \(\omega\) by \(\omega=\dfrac{v}{r}\), so we can solve for \(v\) and have an \(\omega\) to work with!

OpenStudy (theeric):

Do you have any questions? I'm sure you do...

OpenStudy (anonymous):

lol everything well im just stuck because we dont have omega to plug in the second equation and we dont know the velocity either. so idk

OpenStudy (anonymous):

and if we wanted to solve for velocity we dont have omega....

OpenStudy (theeric):

Right! I figured that the whole thing was a bit much! It's complex! I can say more if you need, but let's see if one sentence can clear this up. If we solve for \(v\) in \(\omega=\dfrac{v}{r}\), we can stick it into the other equation. I'm writing more now.

OpenStudy (theeric):

So \(\omega=\dfrac{v}{r}\implies v=\omega r\) \(a_c=\dfrac{v^2}{r}=\dfrac{(\omega r)^2}{r}\) We want \(a_c\), the centripetal acceleration, to be 9.8m/s^2 like gravity. So we know what \(a_c\) is.

OpenStudy (theeric):

We also konw \(r\) since it's given to us, and we want to know \(\omega\).

OpenStudy (anonymous):

AHHHHH

OpenStudy (theeric):

So, after we get \(a_c\), \(r\) and \(\omega\) in one equation, we solve for \(\omega\)! I knew it'd be easier to show than explain :)

OpenStudy (theeric):

Let me know what you get, and let me know if you have any more questions! I suggest that, after you get the answer and I say I agree, you look to see what you actually did so that you can repeat it in the future.

OpenStudy (anonymous):

i got 5.14rads/sec is that right?

OpenStudy (theeric):

That's what I got! Except my calculator said 5.1465, so I would round up to 5.15.

OpenStudy (theeric):

And I agree with your units!

OpenStudy (theeric):

:) Bravo!

OpenStudy (anonymous):

THANK YOU, that only took me forever lol. #lifesaver

OpenStudy (theeric):

Haha, glad I could help! Congrats! You worked through it.